In an extreme test of its braking system under ideal road conditions, a Toyota Celica, travelling initially at 26.9 m/s [S], comes to a stop in 2.61 s. The mass of the car with the driver is 1.18 × 103 kg. Calculate :

(a) the car’s acceleration 

(b) the net force required to cause that acceleration



1
Expert's answer
2022-01-13T09:21:54-0500

a)

"a=\dfrac{v-v_0}{t}=\dfrac{0-26.9\ \dfrac{m}{s}}{2.61\ s}=-10.3\ \dfrac{m}{s^2}."

b)

"F_{net}=ma=1.18\times10^3\ kg\times(-10.3\ \dfrac{m}{s^2})=-12154\ N."